Estimate the values of Zeff for Mg2+ and Al3+.
Answer #1

Answer: i Mght electronic configration = 15225225 Zeft = 2 -o 2= atomic number o = Screening constant Zet (for m 2 ) = 2-6 Z=
Answer #2

9.2 and 8.85